# MagicTailor: Component-Controllable Personalization in Text-to-Image Diffusion Models

### v2

@@ 4 Experimental Results
Figure 6: Qualitative comparisons. We present images generated by MagicTailor and others across various domains. MagicTailor achieves strong text alignment, identity fidelity, and high-quality generation. More results are in the supplementary material.
@@ 4.1 Experimental Setup
For a systematic investigation, we compile a dataset from diverse domains, including characters, animation, buildings, objects, and animals. We use Stable Diffusion 2.1 [35] as the pretrained T2I model and resize all reference images to $512\times 512$. For the warm-up and DS-Bal stages, we set the training steps to 200 and 300, with learning rates of $1\times 10^{-4}$ and $1\times 10^{-5}$, respectively. Leveraging LoRA fine-tuning, each concept-component pair requires only about five minutes of training on an A100 GPU. For evaluation, we design 20 text prompts covering a wide range of scenarios and generate 14,720 images for each method. To ensure fairness, all random seeds are fixed during both training and inference.
@@
We compare our MagicTailor with state-of-the-art (SOTA) personalization methods, including Textual Inversion (TI) [11], DreamBooth (DB) [36], Custom Diffusion (CD) [19], Break-A-Scene (BAS) [2], and CLiC [38]. These methods were selected for their representativeness of personalization frameworks or relevance to learning fine-grained elements. For a fair comparison, we adapt them to our task with minimal modifications, specifically by incorporating the masked diffusion loss (Eq. 1). Apart from method-specific configurations, all methods are implemented using the same setup to ensure consistency. Due to space constraints, additional details on the experimental setup can be found in the supplementary material.
@@ 4.2 Qualitative Comparisons
The qualitative results are shown in in Fig. 6. As observed, TI, CD, and CLiC primarily suffer from semantic pollution, where undesired visual semantics significantly distort the personalized concept. Besides, DB and BAS also struggle in this challenging task, with an overemphasis on either the concept or the component due to semantic imbalance, sometimes even causing the target component to be completely absent. An interesting finding is that imbalanced learning can exacerbate semantic pollution, leading to the color and texture of the target concept or component being mistakenly transferred to unintended parts of the generated images. In contrast, MagicTailor effectively generates text-aligned images that accurately represent both the target concept and component. To further validate its performance, we provide additional comparisons and a generalizability analysis in the supplementary material.
Textual Inversion [11] 0.236 0.742 0.620 0.558
DreamBooth [36] 0.266 0.841 0.798 0.323
Custom Diffusion [19] 0.251 0.797 0.750 0.407
Break-A-Scene [2] 0.259 0.840 0.780 0.338
CLiC [38] 0.263 0.764 0.663 0.499
MagicTailor (Ours) 0.270 0.854 0.813 0.279
Table 2: Quantitative comparisons on the user study. MagicTailor can also outperform other methods across all three aspects in human evaluation.
Textual Inversion [11] 5.8% 2.5% 5.2%
DreamBooth [36] 15.3% 14.7% 12.5%
Custom Diffusion [19] 7.1% 7.7% 9.8%
Break-A-Scene [2] 10.8% 12.1% 22.8%
CLiC [38] 4.5% 5.1% 6.2%
MagicTailor (Ours) 56.5% 57.9% 43.4%
Table 3: Ablation of key techniques. Our DM-Deg and DS-Bal effectively contribute to a superior performance trade-off.
@@ 4.3 Quantitative Comparisons
Automatic Metrics. We utilize four automatic metrics in the aspects of text alignment (CLIP-T [11]) and identity fidelity (CLIP-I [29], DINO [26], DreamSim [10]). To precisely meafsure identity fidelity, we segment out the concept and component in each reference and evaluation image, and then eliminate the target component from the segmented concept. As we can see in Tab. 1, component-controllable personalization remains a tough task even for SOTA methods of personalization. By comparison, MagicTailor can achieve the best results in both identity fidelity and text alignment. It should be credited to the utilization of an effective framework tailored to this special task.
@@
We further evaluate the methods with a user study. Specifically, a detailed questionnaire is designed to display 20 groups of evaluation images with the corresponding text prompt and reference images. Users are asked to select the best result in each group for three aspects, including text alignment, identity fidelity, and generation quality. Finally, we collect a total of 3,180 valid answers and report the selected rates in Tab. 2. It can be observed that MagicTailor can also achieve superior performance in human preferences, showcasing its effectiveness in component-controllable personalization.
Mask-Out Startegy 0.270 0.818 0.760 0.375
Fixed ($\alpha=0.4$) 0.270 0.849 0.800 0.297
Fixed ($\alpha=0.6$) 0.271 0.845 0.794 0.310
Fixed ($\alpha=0.8$) 0.271 0.846 0.796 0.305
Linear (Ascent) 0.270 0.846 0.797 0.307
Linear (Descent) 0.261 0.851 0.802 0.300
Dynamic ($\gamma=8$) 0.266 0.850 0.806 0.289
Dynamic ($\gamma=16$) 0.268 0.854 0.813 0.282
Dynamic ($\gamma=64$) 0.271 0.852 0.812 0.283
Dynamic (Ours) 0.270 0.854 0.813 0.279
4.4 Ablation Studies
We conduct comprehensive ablation studies for MagicTailor to verify the capability of the overall pipeline. More ablation studies are included in the supplementary material.
@@ 4.4 Ablation Studies
In Tab. 3, we investigate two key techniques of MagicTailor by starting from a baseline framework described in Sec. 3.1. Even without DM-Deg and DS-Bal, such a baseline framework can still have competitive performance, showing its reliability. On top of that, we introduce DM-Deg and DS-Bal, where the superior performance trade-off indicates their significance. Qualitative results can also refer to Fig. 2.
Figure 7: Applications of MagicTailor. (a) Decoupled generation. MagicTailor can also separately generate the target concept and component, enriching prospective combinations. (b) Controlling multiple components. MagicTailor shows the potential to handle more than one component, highlighting its effectiveness. (c) Enhancing other generative tools. MagicTailor can seamlessly integrate with various generative tools, adding the capability to control components within their generation pipelines.
Dynamic Intensity Matters.
Momentum Denoising U-Net as a Good Regularizer.
Sensitivity Analysis of Loss Weights.
In Fig. 8, we analyze the sensitivity of loss weights in Eq. 8 (i.e., $\lambda_{\text{pres}}$ and $\lambda_{\text{attn}}$), since loss weights are often critical for model training. As we can see, when $\lambda_{\text{pres}}$ and $\lambda_{\text{attn}}$ vary within a reasonable range, our MagicTailor can consistently attain SOTA performance, revealing the robustness of MagicTailor on these crucial hyperparameters.
Fixed ($\beta=0$) 0.268 0.850 0.803 0.293
Fixed ($\beta=1$) 0.270 0.851 0.808 0.286
Momentum ($\beta=0.5$) 0.268 0.850 0.805 0.290
Momentum ($\beta=0.9$) 0.269 0.850 0.808 0.288
Momentum (Ours) 0.270 0.854 0.813 0.279
4.5 Applications
@@
Figure 8: Ablation of loss weights. We report CLIP-T [11] for text alignment, and DreamSim [10] for identity fidelity as it is most similar to human judgments. For reference, we also present the second-best results in Table 1, highlighting our robustness.
@@
In Fig. 7 (c), we provide some interesting method combinations to show that our MagicTailor can enhance other generative tools. The combined tools include ContorlNet [56], CSGO [51], and InstantMesh [52]. As we can see, MagicTailor can be seamlessly integrated into these tools, furnishing them with an additional ability to control the concept’s component in their pipelines. For instance, working with MagicTailor, InstantMesh can conveniently achieve fine-grained design of 3D mesh, demonstrating the practicability of MagicTailor in collaborative applications.
@@ 5 Conclusion
We introduce component-controllable personalization, enabling precise customization of individual components within a concept. Our framework, MagicTailor, uses Dynamic Masked Degradation (DM-Deg) to suppress unwanted semantics and Dual-Stream Balancing (DS-Bal) to ensure balanced learning. Experiments show that MagicTailor sets a new standard in this task, with promising creative applications. In the future, we would like to extend our approach to broader image and video generation, enabling finer control over multi-level visual semantics for enhanced generative capabilities.
Supplementary Material
@@ A.1 Dataset
As there is no existing dataset specifically for component-controllable personalization, we curate a dataset from the internet to conduct experiments. Particularly, unlike previous works [36, 19] that focus on very few categories of concepts, the dataset contains concepts and components from various domains, such as characters, animation, buildings, objects, and animals. Overall, the dataset consists of 23 concept-component pairs totally with 138 reference images, where each concept/component contains 3 reference images and a corresponding category label. The scale of this dataset is aligned with the scale of those datasets used in the compared methods [11, 36, 19, 2, 38].
@@ A.2 Implementation
We utilize Stable Diffusion 2.1 [35] as the pretrained T2I diffusion model. As commonly done, the resolution of reference images is set to 512 $\times$ 512. Besides, the rank and alpha of the LoRA module are set to 32. To simplify concept learning, we exclude the region of the target component from the segmentation masks of the target concept, e.g., remove the hair from the person in a “<person> + <hair>” pair. For the warm-up and DS-Bal stage, we set the learning rate to 1e-4 and 1e-5 and the training steps to 200 and 300. Moreover, the learning rate is further scaled by the batch size, which is set to completely contain a concept-component pair. For the cross-attention loss, we follow [2] to average the corresponding cross-attention maps at resolution 16 $\times$ 16 and normalized them to [0, 1]. The model is trained with an AdamW [23] optimizer and a DDPM [14] sampler. As done in [2], the tensor precision is set to float16 to accelerate training. For a fair comparison, all random seeds are fixed at 0 in each experiment, and all compared methods use the same implementation above except for method-specific configurations.
@@ A.3 Evaluation
To generate images for evaluation, we carefully design 20 text prompts covering extensive situations, which are listed in Tab. 6. These text prompts can be divided into four aspects, including recontextualization, restylization, interaction, and property modification, where each aspect is composed of 5 text prompts. In recontextualization, we change the contexts to different locations and periods. In restylization, we transfer concepts into various artistic styles. In interaction, we explore the spatial interaction with other concepts. In property modification, we modify the properties of concepts in rendering, views, and materials. Such a group of diverse text prompts allows us to systemically evaluate the generalization capability of a method. We generate 32 images per text prompt for each pair, using a DDIM [41] sampler with 50 steps and a classifier-free guidance scale of 7.5. To ensure fairness, we fix the random seed within the range of [0, 31] across all methods. This process results in a total of 14,720 images for each method to be evaluated, ensuring a robust and thorough comparison.
Table 6: Text prompts used to generate evaluation images. These text prompts can be divided into four aspects: recontextualization, restylization, interaction, and property modification, covering extensive situations to systemically evaluate the generalization capability of a method. Note that “<placeholder>” will be replaced by the combination of pseudo-words (e.g., “<tower> with <roof>”) when generating evaluation images, and will be replaced by the combination of category labels (e.g., “tower with roof”) when calculating the metric of text alignment.
@@ A.4 Automatic Metrics
We utilize four automatic metrics in the aspects of text alignment (CLIP-T [11]) and identity fidelity (CLIP-I [29], DINO [26], DreamSim [10]). To precisely measure identity fidelity, we improve the traditional measurement approach for personalization. This is because a reference image of the target concept/component could contain an undesired component/concept that is not expected to appear in evaluation images. Specifically, we use Grounded-SAM [34] to segment out the concept and component in each reference and evaluation image. Then, we further eliminate the target component from the segmented concept as we have done during training. Such a process is similar to the one adopted in [2]. As a result, using the segmented version of evaluation images and reference images, we can accurately calculate the metrics of identity fidelity.
@@ A.5 User Study
We further evaluate the methods with a user study. Specifically, we design a questionnaire to display 20 groups of evaluation images generated by our method and other methods. Besides, each group also contains the corresponding text prompt and the reference images of the concept and component, where we adopt the same text prompts that are used to calculate CLIP-T. The results of our method and all the compared methods are presented on the same page. Clear rules are established for users to evaluate in three aspects, including text alignment, identity fidelity, and generation quality. Users are requested to select the best result in each group by answering the corresponding questions of these three aspects. We hide all the method names and randomize the order of methods to ensure fairness. Finally, 3,180 valid answers are collected for a sufficient quantitative evaluation of human preferences.

### v3

@@ 4 Experimental Results
Textual Inversion (Gal et al., 2022) 0.236 0.742 0.620 0.558
DreamBooth (Ruiz et al., 2023) 0.266 0.841 0.798 0.323
Custom Diffusion (Kumari et al., 2023) 0.251 0.797 0.750 0.407
Break-A-Scene (Avrahami et al., 2023) 0.259 0.840 0.780 0.338
CLiC (Safaee et al., 2024) 0.263 0.764 0.663 0.499
MagicTailor (Ours) 0.270 0.854 0.813 0.279
Table 2: Quantitative comparisons on the user study. MagicTailor also outperforms other methods in all aspects of human evaluation.
Textual Inversion (Gal et al., 2022) 5.8% 2.5% 5.2%
DreamBooth (Ruiz et al., 2023) 15.3% 14.7% 12.5%
Custom Diffusion (Kumari et al., 2023) 7.1% 7.7% 9.8%
Break-A-Scene (Avrahami et al., 2023) 10.8% 12.1% 22.8%
CLiC (Safaee et al., 2024) 4.5% 5.1% 6.2%
MagicTailor (Ours) 56.5% 57.9% 43.4%
@@ 4.1 Experimental Setup
For a systematic investigation, we collect a dataset from diverse domains, including characters, animation, buildings, objects, and animals. We use Stable Diffusion (SD) 2.1 (Rombach et al., 2022) as the pretrained T2I model. For the warm-up and DS-Bal stages, we set the training steps to 200 and 300, with learning rates of $1\times 10^{-4}$ and $1\times 10^{-5}$, respectively. Each concept-component pair requires only about five minutes of training on an A100 GPU. For evaluation, we design 20 text prompts covering a wide range of scenarios and generate 14,720 images for each method. To ensure fairness, all random seeds are fixed during both training and inference. More details of the experimental setup are included in Appendix A.
@@
We compare our MagicTailor with several personalization methods, including Textual Inversion (TI) (Gal et al., 2022), DreamBooth (DB) (Ruiz et al., 2023), Custom Diffusion (CD) (Kumari et al., 2023), Break-A-Scene (BAS) (Avrahami et al., 2023), and CLiC (Safaee et al., 2024). These methods were selected for their representativeness of personalization frameworks or relevance to learning fine-grained elements. For a fair comparison, we adapt them to our task with minimal modifications, specifically by incorporating the masked diffusion loss (Eq. 1). Apart from method-specific configurations, all methods are implemented using the same setup to ensure consistency.
@@ 4.2 Qualitative Comparisons
The qualitative results are shown in in Fig. 6. As observed, TI, CD, and CLiC primarily suffer from semantic pollution, where undesired visual semantics significantly distort the personalized concept. Besides, DB and BAS also struggle in this challenging task, with an overemphasis on either the concept or the component due to semantic imbalance, sometimes even causing the target component to be completely absent. An interesting finding is that imbalanced learning can exacerbate semantic pollution, leading to the color and texture of the target concept or component being mistakenly transferred to unintended parts of the generated images. In contrast, MagicTailor effectively generates text-aligned images that accurately represent both the target concept and component. To further demonstrate the performance of MagicTailor, we provide additional comparisons in Appendix B.
Table 3: Effectiveness of key techniques. Our DM-Deg and DS-Bal effectively contribute to a superior performance trade-off.
@@
Figure 7: Compatibility with different backbones. We equip MagicTailor with SD 1.5 (Rombach et al., 2022), SD 2.1 (Rombach et al., 2022), and SDXL (Podell et al., 2023). The results show that MagicTailor can be generalized to multiple backbones, and a better backbone could provide better generation quality.
Figure 8: Robustness on loss weights. We report CLIP-T (Gal et al., 2022) for text alignment, and DreamSim (Fu et al., 2023) for identity fidelity as it is most similar to human judgments. Second-best results in Table 1 are also presented to highlight our robustness.
Figure 9: Performance on different numbers of reference images. We present qualitative results to show that MagicTailor can still achieve satisfactory performance when provided only 1 or 2 reference image(s) per concept and component.c
@@ 4.3 Quantitative Comparisons
Automatic Metrics. We utilize four automatic metrics in the aspects of text alignment (CLIP-T (Gal et al., 2022)) and identity fidelity (CLIP-I (Radford et al., 2021), DINO (Oquab et al., 2023), DreamSim (Fu et al., 2023)). To precisely measure identity fidelity, we segment out the concept and component in each reference and evaluation image, and then eliminate the target component from the segmented concept. As we can see in Tab. 1, component-controllable personalization remains a tough task even for SOTA methods of personalization. By comparison, MagicTailor achieves the best results in both identity fidelity and text alignment. It should be credited to the effective framework tailored to this special task.
@@
We further evaluate the methods with a user study. Specifically, a detailed questionnaire is designed to display 20 groups of evaluation images with the corresponding text prompt and reference images. Users are asked to select the best result in each group for three aspects, including text alignment, identity fidelity, and generation quality. Finally, we collect a total of 3,180 valid answers and report the selected rates in Tab. 2. It can be observed that MagicTailor can also achieve superior performance in human preferences, further verifying its effectiveness.
4.4 Ablation Studies and Analysis
We conduct comprehensive ablation studies and analysis for MagicTailor to verify its capability. More ablation studies and analysis are included in Appendix C.
@@ 4.4 Ablation Studies and Analysis
In Tab. 3, we investigate two key techniques by starting from a baseline framework described in Sec. 3.1. Even without DM-Deg and DS-Bal, such a baseline framework can still have competitive performance, showing its reliability. On top of that, we introduce DM-Deg and DS-Bal, where the superior performance trade-off indicates their significance. Qualitative results can refer to Fig. 2.
Figure 10: Further applications of MagicTailor. (a) Decoupled generation: MagicTailor can also separately generate the target concept and component, enriching prospective combinations. (b) Controlling multiple components: MagicTailor shows the potential to handle more than one component, highlighting its effectiveness. (c) Enhancing other generative tools: MagicTailor can seamlessly integrate with various generative tools, adding the capability to control components within their generation pipelines.
Figure 11: Generalizability for complex prompts. We present qualitative results generated with complex text prompts. In addition to those well-categorized text prompts, our MagicTailor can also follow more complex ones to generate text-aligned images.
Figure 12: Generalizability for difficult pairs. We show the results of two hard cases involving large geometric discrepancy and cross-domain interactions, showing that MagicTailor can effectively handle such challenging scenarios.
Compatibility with Different Backbones.
MagicTailor can also collaborate with other T2I diffusion models as it is a model-independent approach. In Fig. 7, we employ MagicTailor in other backbones like SD 1.5 (Rombach et al., 2022) and SDXL (Podell et al., 2023), showcasing MagicTailor can also achieve remarkable results. Notably, we directly use the original hyperparameter values without further selections, showing the generalizability of MagicTailor.
Robustness on Loss Weights.
In Fig. 8, we analyze the sensitivity of loss weights in Eq. 8 (i.e., $\lambda_{\text{pres}}$ and $\lambda_{\text{attn}}$), since loss weights are often critical for model training. As we can see, when $\lambda_{\text{pres}}$ and $\lambda_{\text{attn}}$ vary within a reasonable range, our MagicTailor can consistently attain SOTA performance, revealing its robustness on these hyperparameters.
Performance on Different Numbers of Reference Images.
In Fig. 9, we reduce the number of reference images to analyze the performance variation. With fewer reference images, MagicTailor can still show satisfactory results. While more reference images could lead to better generalization ability, one reference image per concept/component is enough to obtain a decent result with our MagicTailor.
Generalizability to Complex Prompts.
In comparisons, we have used well-categorized text prompts for systemic evaluation. Here we further evaluate MagicTailor’s performance on other complex text prompts involving more complicated contexts. As shown in Fig. 11, MagicTailor effectively generates text-aligned images when performing fidelity personalization, showing its ability to handle diverse user needs.
Generalizability to Difficult Pairs.
We further evaluate MagicTailor’s performance on challenging pairs, focusing on two cases: 1) large geometric discrepancy, such as “<person>” in an upper body portrait and “<hair>” in a profile photo, and 2) cross-domain interactions, such as “<person>” and “<ear>” of dogs. As shown in Fig. 12, even facing these hard cases, MagicTailor can still effectively personalize target concepts and components with high fidelity.
4.5 Further Applications
@@
We demonstrate how MagicTailor enhances other generative tools like ControlNet (Zhang et al., 2023), CSGO (Xing et al., 2024), and InstantMesh (Xu et al., 2024) in Fig. 10(c). MagicTailor can integrates seamlessly, furnishing them with an additional ability to control the concept’s component in their pipelines. For instance, working with MagicTailor, InstantMesh can conveniently achieve fine-grained 3D mesh design, exhibiting the practicability of MagicTailor in more creative applications.
@@ 5 Conclusion
We introduce component-controllable personalization, enabling precise customization of individual components within concepts. The proposed MagicTailor uses Dynamic Masked Degradation (DM-Deg) to suppress unwanted semantics and Dual-Stream Balancing (DS-Bal) to ensure balanced learning. Experiments show that MagicTailor sets a new standard in this task, with promising creative applications. In the future, we would like to extend our approach to broader image and video generation, enabling finer control over multi-level visual semantics for creative generation capabilities.
Acknowledgments
We would like to thank Pengzhi Li, Tian Ye, Jinyu Lin, and Jialin Gao for their valuable discussion and suggestions. This study was supported by the InnoHK initiative of the Innovation and Technology Commission of the Hong Kong Special Administrative Region Government via the Hong Kong Centre for Logistics Robotics.
@@ A.1 Dataset
As there is no existing dataset specifically for component-controllable personalization, we curate a dataset from the internet to conduct experiments. Particularly, unlike previous works [Ruiz et al., 2023; Kumari et al., 2023] that focus on very few categories of concepts, the dataset contains concepts and components from various domains, such as characters, animation, buildings, objects, and animals. Overall, the dataset consists of 23 concept-component pairs totally with 138 reference images, where each concept/component contains 3 reference images and a corresponding category label. It is worth noting that the scale of this dataset is aligned with the scale of those datasets used in the compared methods [Gal et al., 2022; Ruiz et al., 2023; Kumari et al., 2023; Avrahami et al., 2023; Safaee et al., 2024].
@@ A.2 Implementation
We utilize SD 2.1 [Rombach et al., 2022] as the pretrained T2I diffusion model. As commonly done, the resolution of reference images is set to 512 $\times$ 512. Besides, the LoRA rank and alpha are set to 32. To simplify concept learning, we exclude the region of the target component from the segmentation masks of the target concept, e.g., remove the hair from the person in a “<person> + <hair>” pair. For the warm-up and DS-Bal stage, we set the learning rate to 1e-4 and 1e-5 and the training steps to 200 and 300. Moreover, the learning rate is further scaled by the batch size, which is set to completely contain a concept-component pair. For the cross-attention loss, we follow [Avrahami et al., 2023] to average the corresponding cross-attention maps at resolution 16 $\times$ 16 and normalized them to [0, 1]. The model is trained with an AdamW [Loshchilov and Hutter, 2017] optimizer and a DDPM [Ho et al., 2020] sampler. As done in [Avrahami et al., 2023], the tensor precision is set to float16 to accelerate training. For a fair comparison, all random seeds are fixed at 0, and all compared methods use the same implementation above except for method-specific configurations.
@@ A.3 Evaluation
To generate images for evaluation, we carefully design 20 text prompts covering extensive situations, which are listed in Tab. 4. These text prompts can be divided into four aspects, including recontextualization, restylization, interaction, and property modification, where each aspect is composed of 5 text prompts. In recontextualization, we change the contexts to different locations and periods. In restylization, we transfer concepts into various artistic styles. In interaction, we explore the spatial interaction with other concepts. In property modification, we modify the properties of concepts in rendering, views, and materials. Such a group of diverse text prompts allows us to systemically evaluate the generalization capability of a method. We generate 32 images per text prompt for each pair, using a DDIM [Song et al., 2020] sampler with 50 steps and a classifier-free guidance scale of 7.5. To ensure fairness, we fix the random seed within the range of [0, 31] across all methods. This process results in 14,720 images for each method to be evaluated, ensuring a thorough comparison.
Table 4: Text prompts used to generate evaluation images. These text prompts can be divided into four aspects: recontextualization, restylization, interaction, and property modification, covering extensive situations to systemically evaluate the method’s generalizability. Note that “<placeholder>” will be replaced by the combination of pseudo-words (e.g., “<tower> with <roof>”) when generating evaluation images, and will be replaced by the combination of category labels (e.g., “tower with roof”) when calculating the metric of text alignment.
@@ A.4 Automatic Metrics
We utilize four automatic metrics in the aspects of text alignment (CLIP-T [Gal et al., 2022]) and identity fidelity (CLIP-I [Radford et al., 2021], DINO [Oquab et al., 2023], DreamSim [Fu et al., 2023]). To precisely measure identity fidelity, we improve the traditional measurement approach for personalization. This is because a reference image of the target concept/component could contain an undesired component/concept that is not expected to appear in evaluation images. Specifically, we use Grounded-SAM [Ren et al., 2024] to segment out the concept and component in each reference and evaluation image. Then, we further eliminate the target component from the segmented concept as we have done during training. Such a process is similar to the one adopted in [Avrahami et al., 2023]. As a result, using the segmented version of evaluation images and reference images, we can accurately calculate the metrics of identity fidelity.
@@ A.5 User Study
We further evaluate the methods with a user study. Specifically, we design a questionnaire to display 20 groups of evaluation images generated by our method and other methods. Besides, each group also contains the corresponding text prompt and the reference images of the concept and component, where we adopt the same text prompts that are used to calculate CLIP-T. The results of our method and all the compared methods are presented on the same page. Clear rules are established for users to evaluate in three aspects, including text alignment, identity fidelity, and generation quality. Users are requested to select the best result in each group by answering the corresponding questions of these three aspects. We hide all the method names and randomize the order of methods to ensure fairness. Finally, 3,180 valid answers are collected for a sufficient evaluation of human preferences.